OpenStudy (j.f):

Metal f^-1(x) Show the steps it takes to arrive at the answer (x-1)/2

OpenStudy (xapproachesinfinity):

and where f(x)? you can't do f inverse just like that, need more info

OpenStudy (j.f):

I'm sorry. f(x)= 2x+1 and g(x)= x^2-1

OpenStudy (xapproachesinfinity):

i will coach you to do it and you do it for g(x)

OpenStudy (xapproachesinfinity):

f(x)=2x+1 so first do y=2x+1

OpenStudy (xapproachesinfinity):

okay

OpenStudy (xapproachesinfinity):

hey say yes no! why! i need to see you interact

OpenStudy (xapproachesinfinity):

i don't like to just speak with myself here!

OpenStudy (xapproachesinfinity):

you are not even caring, while i was replying you are checking around hehe

OpenStudy (j.f):

so y=2x+1=2y=x+1= y=(x+1)/2. How does the 1 become a negative?